solve for x. 2/3 (x+9) = 1/3 (12-x)
Answer by deborabr(181)

2/3 (x+9) = 1/3 (12-x)
2(x+9)/3 = (12-x)/3
(2x+18)/3 = (12-x)/3
with equal denominators we can delete them
2x+18=12-x --> x with x and single numbers together too
2x+x=12-18 --> calculate them
3x=-6 --> pass the number that is multiplying dividing
x=-6/3 and calculate
x=-2